OpenStudy (anonymous):

a person wishes to invest M dollars at t eh end of each month from January 2000 until the end of December 2003. If the account gives interest at the annual rate of 18% compounded monthly and the individual wishes to have $100,000 by the end of 2003, how much should be invested each month?

OpenStudy (anonymous):

No i don't

OpenStudy (anonymous):

it is \[A=P\times{(1+r)^{n}}\] where A = Amount P = Principal r = Rate of interest n = number of time period

OpenStudy (anonymous):

ok now tell me for how many moths you will be investing??

OpenStudy (anonymous):

so for 3 years that's 36 months

OpenStudy (anonymous):

you started investing from Jan 200 - Dec 2003 right it is 36 months

OpenStudy (anonymous):

A=100,000x(1+18%)^36

OpenStudy (anonymous):

a=100,000 (19)^36

OpenStudy (anonymous):

so here n = 3yrs = 36 months r = 18% per annum but this will get compounded monthly so it cannot be 18%

OpenStudy (anonymous):

can i leave it in the annual form and just do A=100,000x(19)^3 ?

OpenStudy (anonymous):

so your formula will change right the formula that i had given u is the general formula now when it is compounded monthly in that case the formula would be \[A = P\times{(1+{{r}\over{12}})^{12n}}\]

OpenStudy (anonymous):

@chrissytt17 here principal is unknown

OpenStudy (anonymous):

are you getting my point?

OpenStudy (anonymous):

i thought the principal is the amount that the individual wants

OpenStudy (anonymous):

no what you want that you will get at what time????

OpenStudy (anonymous):

the individual wishes to have $100,000 by the end of 2003, how much should be invested each month?

OpenStudy (anonymous):

100,000=Px(1+r/12)^12n

OpenStudy (anonymous):

where r = 18% but when you are plugging in the value of r that time convert it into decimal so r = 0.18

OpenStudy (anonymous):

100,000= P x (1+.18/12)^12*3 ?

OpenStudy (anonymous):

100,000= P x (1.015)^36

OpenStudy (anonymous):

am i on the right track?

OpenStudy (anonymous):

wait let me cross check it

OpenStudy (anonymous):

i came up with about 58,513.75

OpenStudy (anonymous):

no it is not correct

OpenStudy (anonymous):

I followed the formula

OpenStudy (anonymous):

I think I am lost

OpenStudy (anonymous):

hey i was not telling you about the Annuity do you know about it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you are in which class

OpenStudy (anonymous):

math 118

OpenStudy (anonymous):

ok do you know about time value of money

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

\[100000 = P \times \sum_{n=1}^{36}(1.015)^n\]\]

OpenStudy (anonymous):

ok solve this one you will get the exact value

OpenStudy (anonymous):

the second factor on RHS is called as Present Value Interest Factor of Annuity

OpenStudy (anonymous):

I don't know what that symbol is

OpenStudy (anonymous):

it is summation